# LaRa: Efficient Large-Baseline Radiance Fields

### v1

@@ 1 Introduction
Despite these successes, geometry with view-consistent textures is not the only aspect required in applications of 3D reconstruction. The reconstruction process should also be able to recover view-dependent appearance. To this end, neural radiance fields [42] and neural implicit surfaces [47, 74] investigate volumetric representations that can be learned from multi-view captures without explicit feature matching. Their follow-ups [64, 76, 44, 21, 5, 32, 57, 2, 77] improve efficiency and quality, but mostly require per-scene optimization and dense multi-view supervision.
@@
In this work, we present LaRa, a feed-forward reconstruction model without the requirement of heavy training resources for the task of 360 bounded radiance fields reconstruction from unstructured few-views. The core idea of our work is to progressively and implicitly perform feature matching through a novel volume transformer. We propose a Gaussian volume as 3D representation, in which each voxel comprises a set of learnable Gaussian primitives. To obtain the Gaussian volume from image conditions, we progressively update a learnable embedding volume by querying features in 3D. Specifically, we utilize a DINO image feature encoder to obtain image tokens and lift 2D tokens to 3D by unprojecting them to a shared canonical space. Next, we propose a novel Group Attention Layer architecture to enable local and global feature aggregation. Specifically, we divide dense volumes into local groups and only apply attention within each group, inspired by standard feature point matching. The grouped features and embeddings are fed to a cross-attention sub-layer to implicitly match features between feature groups of the feature volume and embedding volume, which is followed by a 3D CNN layer to efficiently share information across neighboring groups. After passing through all attention layers, the volume transformer outputs a Gaussian volume, and is then decoded as 2D Gaussian [27] parameters using a coarse-to-fine decoding process. By incorporating efficient rasterization, our method achieves high-resolution renderings.
@@ 2 Related Work
Multi-view stereo.   Multi-view stereo reconstruction aims to generate detailed 3D models by reasoning from images captured from multiple viewpoints, which has been studied for decades [14, 35, 33, 25, 52, 22, 50]. In recent years, multi-view stereo networks [72, 28] have been proposed to address MVS problems. MVSNet [72] utilizes a 3D Convolutional Neural Network for processing a cost volume. This cost volume is created by aggregating features from a set of adjacent views, employing the plane-sweeping technique from a reference viewpoint, facilitating depth estimation and enabling superior 3D reconstructions. Subsequent research has built on top of this foundation, incorporating strategies such as iterative plane sweeping [73], point cloud enhancement [9], confidence-driven fusion [41], and the usage of multiple cost volumes [12, 24] to further refine reconstruction accuracy. However, all of these works require large image overlap for faithful feature matching.
@@
Recently, several works have explored feed-forward models for few-shot [1, 36, 19, 43, 4, 11, 67] input by capitalizing on large-scale training datasets and model sizes. They leverage cross-view attention to globally reason about 3D scenes and output 3D representation (e.g., tri-plane, IB-planes) for radiance field reconstruction. Concurrent work by LGM [59] and GRM [75] introduces few-shot 3D reconstruction models that produce high-resolution 3D Gaussians using a transformer framework. While these methods achieve impressive visual results, training becomes expensive and less practical for the academic community. Unlike some recent single view reconstruction methods [58, 26], our work focuses on few-shot ($>1$) reconstruction since single-view input can be efficiently lifted to multi-view by multi-view generative models [38, 71, 54, 39].
@@ 3 LaRa: Large-baseline Radiance Fields
Our goal is to reconstruct the geometry and view-dependent appearance of bounded scenes from sparse input views using limited training resources. Given $M$ images $\mathbf{I}\!=\!(I_{1},\ldots,I_{M})$ with camera parameters $\boldsymbol{\pi}\!=\!(\pi_{1},\ldots,\pi_{M})$, our method reconstructs radiance fields as a collection of 2D Gaussians, which is used to synthesize novel views and extract meshes. Our model is a function $\mathbf{f}$ of a discrete radiance field of voxel positions $\mathbf{v}$ and outputs a Gaussian volume $\mathbf{V}_{\mathcal{G}}$
@@
Image feature volume.   We construct a feature volume for each input view by lifting the 2D image features to a canonical volume defined in the center of the scene. We use the DINO [3] image encoder to extract per-view image features, and inject Plücker ray directions into the features via adaptive layer norm [48]. Unlike previous works that modulate camera poses to image features using extrinsic and intrinsic matrices [38, 26], Plücker rays are defined by the cross product between the camera location and ray direction, offering a unique ray parameterization independent of object scale, camera position and focal length. After modulation, we obtain $M$ per-view image feature maps. We further lift the 2D maps to 3D by back-projecting the feature maps to a canonical volume, therefore resulting in $M$ feature volumes $\mathbf{V}_{\text{f}}\in\mathbb{R}^{W\times W\times W\times O}$ with $O$ channels.
Embedding volume.   Inspired by prior works [31, 45, 26], we construct a learnable embedding volume $\mathbf{V}_{\text{e}}\in\mathbb{R}^{W\times W\times W\times C}$ for modeling prior knowledge. 3D reconstruction is generally under-constrained in sparse view settings, hence prior knowledge is critical for faithful reconstructions. We propose to leverage a 3D embedding volume to model and learn prior information across objects, which acts as a 3D object template that greatly reduces the solution space. The embedding volume is aligned with the image feature volume, allowing for efficient cross attention (see Section 3.2).
@@ 3.2 Volume Transformer
We now present our novel volume transformer containing a set of group attention layers that progressively update the embedding volume. Our group attention layers contain three sublayers (see Figure 2): group cross-attention, a multi-layer perceptron (MLP), and 3D convolution. Given the image feature volume and embedding volume, we first unfold these 3D volumes (i.e., $\mathbf{V}_{\text{f}}$ and $\mathbf{V}_{\text{e}}$) into $G$ local token groups along each axis. We then apply a cross-attention layer between the corresponding groups of embedding tokens $\mathbf{V}_{\text{e}}^{g,j}$ and image feature tokens $\mathbf{V}^{g}_{\text{f}}$, where $j$ denotes the index of the layer starting from 1, and $\{\mathbf{V}_{\text{e}}^{g,1}\}_{g}\!=\!\mathbf{V}_{\text{e}}$. Figure 1 illustrates the unfolding for $G=4$ and highlights the corresponding groups.
@@ 3.3 Coarse-Fine Decoding
We obtain 2D Gaussian primitive shape and appearance parameters from the Gaussian volume, so we introduce a coarse-fine decoding process to better recover texture details. Instead of using a single network and sampling scheme to reason about the scene, we simultaneously optimize two decoding modules: one “coarse” and one “fine”.
@@
Despite the fact that the coarse renderings can already provide faithful depths/geometries, the appearance tends to be blurred, as shown in (e) of  Figure 6. This is because the image texture can easily be lost after the DINO encoder and the Group Attention layers. To address this problem, we propose a “fine” decoding module to guide fine texture prediction.
@@
In practice, the depth features can change significantly across different scenes. To mitigate scaling discrepancies, we replace the rendering depth $\hat{\mathbf{D}}_{\mathbf{p}^{k}_{i}}$ with a displacement feature $\left\lvert\hat{\mathbf{D}}_{\mathbf{p}^{k}_{i}}-z_{\mathbf{p}^{k}_{i}}\right\rvert$ that compares the rendered depth for input views and the depth $z_{\mathbf{p}^{k}_{i}}$ of a primitive, allowing for occlusion-aware reasoning.
@@
$\text{SH}_{i,k}^{\textit{residuals}}$ $=\text{MLP}\left(\text{CrossAttn}\left(\mathcal{X}_{\mathbf{p}^{k}_{i}},\mathbf{V}^{i}_{\text{e}}\right)\right)\text{,}$ (6)
@@
Intuitively, the “fine” decoding module attempts to learn a geometry-aware texture blending process based on multi-view images, primitive features, and rendering buffers from the coarse module. Furthermore, both coarse and fine modules are differentiable and updated simultaneously. Thus, the fine renderings can further regularize the coarse predictions.
@@ 4 Implementation Details
Datasets.   We train our model on multi-view synthetic renderings of objects [69], based on the Objaverse dataset [15], which includes 264,775 scenes with a train/test split of 10:1. Each scene contains 38 circular views with an image resolution of $512\times 512$. To ensure sufficient angular coverage of the input views, we employ the classical K-means algorithm to cluster the cameras into 4 clusters. During training, we randomly choose two views from each cluster for every iteration, in which the first 4 images share the same camera poses as the input views, while the remaining 4 images are novel view outputs. We employ the eight output images for supervision and leverage the loss objectives outlined in Eq. 8 to update the network.
@@
Network.   We developed LaRa using PyTorch Lightning [20] and conduct our training on 4 NVIDIA A100-40G GPUs over a period of 2 days for the fast model and 3.5 days for the base model, with a batch size of 2 per GPU. We use DINO-base for encoding $M\!=\!4$ multi-view images at a resolution of $512\times 512$. We use a volume resolution of $W\!=\!16$ with $C\!=\!768$ channels for the image feature volume, and a resolution of $W\!=\!32$ with $C\!=\!256$ channels for the embedding volume, dividing both into $G\!=\!16$ groups for the group attention layers. Our group attention network consists of 12 layers, producing a Gaussian volume of size $64\!\times\!64\!\times\!64\!\times\!80$. We choose $K\!=\!2$ primitives for each voxel, and constrain the offset radius to $r=1/32$ in our experiments. The total number of trainable parameters is 125 million.
@@ 5.1 Comparison
Our approach is robust to scene scale and can generalize to real captured images, such as those in the Co3D dataset, thanks to our canonical modeling and projection-based feature lifting. In contrast, LGM [59] leverages a monocular prediction and fusion technique that requires a reference scene scale and a constant camera-object distance to avoid focal length and distance ambiguity. This requirement significantly limits its generalizability to real data. As shown in Table 1 and  Figure 4, LGM provides faithful reconstructions in datasets with a strict constant camera-object distance, such as GSO, but fails to generalize to unconstraint multi-view data such as in Objaverse and Co3D datasets, and exhibits serious distortions. Our model trained on 4 A100-40G GPUs for 2 days demonstrates superior results compared to the LGM model trained on 32 A100-80G GPUs (8$\times$ GPUs, 16$\times$ RAM, 32$\times$ GPU hours) and on the same synthetic Objaverse dataset [15].
@@
Geometry.   We evaluate the quality of our geometry reconstruction by comparing the depth reconstructions on novel views, generated by a weighted sum of the $z$ values of the primitives. As shown in Table 2, our approach achieves significantly lower L1 errors and higher geometry accuracy other baselines. In Figure 5, we also visualize geometry reconstruction by extracting meshes using TSDF. In addition, our trajectory video rendering (48 views at a resolution of $512$) together with mesh extraction is highly efficient, as it does not require fine-tuning and can be performed in just 2 seconds.
@@ 5.2 Ablation Study
We now analyze the contributions of individual elements of our model design. To reduce the training cost, we reduce the training from 50 to 30 epochs for ablations.
@@
Effect of input views.   Our approach is highly efficient and compatible with different numbers of input views. In prior experiments, we utilize 4 views for both training and inference as our standard configuration. We evaluate our method in $2-4$ testing views (as shown in rows (h),(j), and (g)) using the full model in row (g).
@@ 6 Conclusion
We have presented LaRa, a novel method for 360 bounded radiance fields reconstruction from large-baseline inputs. Our central idea is to match image features and embedding volume through unified local and global attention layers. By integrating this with a coarse-fine decoding and splatting process, we achieve high efficiency for both training and inference. In future work, we plan to explore how to enlarge the batch size per-GPU and volume resolution without increasing GPU usage. In addition, we hope to investigate how to extend it to handle unbounded 360 scenes and decompose the radiance field into its constituent physical material and lighting components.
Limitations and Discussions.   Our LaRa demonstrates a remarkable efficiency feed-forward model that achieved high-fidelity all-around novel-view synthesis and surface reconstruction from sparse large-baseline images. However, our approach struggles to recover high-frequency geometry and texture details, mainly due to the low volume resolution. Enhancing our approach with techniques such as gradient checkpointing or mixed-precision training can potentially increase training batch size as well as volume resolution. We have also noticed that our method can yield inconsistent rendering results when the geometry is incorrectly estimated or when reconstructing multi-view inconsistent inputs, as demonstrated in the comparison video. This occurs because our method utilizes second-order Spherical Harmonic appearance modeling. While such modeling can capture view-dependent effects, it also introduces a stronger ambiguity between geometry and appearance. We believe that incorporating our method with a physically-based rendering process can potentially address this issue. In addition, our work assumes posed inputs, but estimating precise camera poses for sparse views is a challenge in practice. Incorporating a pose estimation module [65] into the feed-forward setting is an orthogonal direction to our work.
7 Acknowledgements
We thank Bozidar Antic for pointing out a bug that resulted in an improvement of about 1dB. Special thanks to BinBin Huang and Zehao Yu for their helpful discussion and suggestions. We would like to thank Bi Sai, Jiahao Li, Zexiang Xu for providing us with the testing examples of Instant3D, and Jiaxiang Tang for helping us to construct a comparison with LGM.

### v2

@@ Abstract
Figure 1: LaRa is a feed-forward 2D Gaussian Splatting model that reconstructs radiance fields from large-baseline views, a single image, or a text prompt.
@@ 1 Introduction
Despite these successes, geometry with view-consistent textures is not the only aspect required in applications of 3D reconstruction. The reconstruction process should also be able to recover view-dependent appearance. To this end, neural radiance fields [42] and neural implicit surfaces [47, 75] investigate volumetric representations that can be learned from multi-view captures without explicit feature matching. Their follow-ups [65, 77, 44, 21, 5, 32, 58, 2, 78] improve efficiency and quality, but mainly require per-scene optimization and dense multi-view supervision.
@@
In this work, we present LaRa, a feed-forward reconstruction model without the requirement of heavy training resources for the task of 360 bounded radiance fields reconstruction from unstructured few-views. The core idea of our work is to progressively and implicitly perform feature matching through a novel volume transformer. We propose a Gaussian volume as the 3D representation, in which each voxel comprises a set of learnable Gaussian primitives. To obtain the Gaussian volume from image conditions, we progressively update a learnable embedding volume by querying features in 3D. Specifically, we utilize a DINO image feature encoder to obtain image tokens and lift 2D tokens to 3D by unprojecting them into a shared canonical space. Next, we propose a novel Group Attention Layer architecture to enable local and global feature aggregation. Specifically, we divide dense volumes into local groups and only apply attention within each group, inspired by standard feature point matching. The grouped features and embeddings are fed to a cross-attention sub-layer to implicitly match features between feature groups of the feature volume and embedding volume, which is followed by a 3D CNN layer to efficiently share information across neighboring groups. After passing through all attention layers, the volume transformer outputs a Gaussian volume, and is then decoded as 2D Gaussian [27] parameters using a coarse-to-fine decoding process. By incorporating efficient rasterization, our method achieves high-resolution renderings.
@@ 2 Related Work
Multi-view stereo.   Multi-view stereo reconstruction aims to generate detailed 3D models by reasoning from images captured from multiple viewpoints, which has been studied for decades [14, 35, 33, 25, 53, 22, 51]. In recent years, multi-view stereo networks [73, 28] have been proposed to address MVS problems. MVSNet [73] utilizes a 3D Convolutional Neural Network for processing a cost volume. This cost volume is created by aggregating features from a set of adjacent views, employing the plane-sweeping technique from a reference viewpoint, facilitating depth estimation and enabling superior 3D reconstructions. Subsequent research has built on top of this foundation, incorporating strategies such as iterative plane sweeping [74], point cloud enhancement [9], confidence-driven fusion [41], and the usage of multiple cost volumes [12, 24] to further refine reconstruction accuracy. However, all of these works require a large image overlap to achieve faithful feature matching.
@@
Recently, several works have explored feed-forward models for few-shot [1, 36, 19, 43, 4, 11, 68] input by capitalizing on large-scale training datasets and model sizes. They leverage cross-view attention to global reason about 3D scenes and output 3D representation (e.g., tri-plane, IB-planes) for radiance field reconstruction. Concurrent work by LGM [60] and GRM [76] introduces few-shot 3D reconstruction models that produce high-resolution 3D Gaussians using a transformer framework. While these methods achieve impressive visual results, training becomes expensive and less practical for the academic community. Unlike some recent single view reconstruction methods [59, 26], our work focuses on few-shot ($>1$) reconstruction since single-view input can be efficiently lifted to multi-view by multi-view generative models [38, 72, 55, 39].
@@ 3 LaRa: Large-baseline Radiance Fields
Our goal is to reconstruct the geometry and view-dependent appearance of bounded scenes from sparse input views using limited training resources. Given $M$ images $\mathbf{I}\!=\!(I_{1},\ldots,I_{M})$ with camera parameters $\boldsymbol{\pi}\!=\!(\pi_{1},\ldots,\pi_{M})$, our method reconstructs radiance fields as a collection of 2D Gaussians, which is used to synthesize novel views and extract meshes. Our model is a function $\mathbf{f}$ of a discrete radiance field of voxel positions $\mathbf{v}$ and produces a Gaussian volume $\mathbf{V}_{\mathcal{G}}$
@@
Image feature volume.   We construct a feature volume for each input view by lifting the 2D image features to a canonical volume defined in the center of the scene. We use the DINO [3] image encoder to extract per-view image features, and inject Plücker ray directions into the features via the adaptive layer norm [48]. Unlike previous works that modulate camera poses to image features using extrinsic and intrinsic matrices [38, 26], Plücker rays are defined by the cross product between the camera location and ray direction, offering a unique ray parameterization independent of object scale, camera position, and focal length. After modulation, we obtain $M$ per-view image feature maps. We further lift the 2D maps to 3D by back-projecting the feature maps to a canonical volume, therefore resulting in $M$ feature volumes $\mathbf{V}_{\text{f}}\in\mathbb{R}^{W\times W\times W\times O}$ with $O$ channels.
Embedding volume.   Inspired by prior works [31, 45, 26], we construct a learnable embedding volume $\mathbf{V}_{\text{e}}\in\mathbb{R}^{W\times W\times W\times C}$ to model prior knowledge. 3D reconstruction is generally under-constrained in sparse view settings, hence prior knowledge is critical for faithful reconstructions. We propose to leverage a 3D embedding volume to model and learn prior information across objects, which acts as a 3D object template that greatly reduces the solution space. The embedding volume is aligned with the image feature volume, allowing for efficient cross attention (see Section 3.2).
@@ 3.2 Volume Transformer
We now present our novel volume transformer containing a set of group attention layers that progressively update the embedding volume. Our group attention layers contain three sublayers (see Figure 3): group cross-attention, a multi-layer perceptron (MLP), and 3D convolution. Given the image feature volume and embedding volume, we first unfold these 3D volumes (i.e., $\mathbf{V}_{\text{f}}$ and $\mathbf{V}_{\text{e}}$) into $G$ groups along each axis. We then apply a cross-attention layer between the corresponding groups of embedding tokens $\mathbf{V}_{\text{e}}^{g,j}$ and image feature tokens $\mathbf{V}^{g}_{\text{f}}$, where $j$ denotes the index of the layer starting from 1, and $\{\mathbf{V}_{\text{e}}^{g,1}\}_{g}\!=\!\mathbf{V}_{\text{e}}$. Figure 2 illustrates the unfolding of $G=4$ and highlights the corresponding groups.
@@ 3.3 Coarse-Fine Decoding
We obtain 2D Gaussian primitive shape and appearance parameters from the Gaussian volume, so we introduce a coarse-fine decoding process to better recover texture details. Instead of using a single network and sampling scheme to reason about the scene, we simultaneously optimize two decoding modules: one “coarse” and one “fine”. Intuitively, the “fine” decoding module attempts to learn a geometry-aware texture blending process based on multi-view images, primitive features, and rendering buffers from the coarse module.
@@
Despite the fact that the coarse renderings can already provide faithful depths/geometries, the appearance tends to be blurred, as shown in (e) of  Figure 7. This is because the image texture can easily be lost after the DINO encoder and the Group Attention layers. To address this problem, we propose a “fine” decoding module to guide the fine texture prediction.
@@
In practice, the depth features can change significantly in different scenes. To mitigate scaling discrepancies, we replace the rendering depth $\hat{\mathbf{D}}_{\mathbf{p}^{k}_{i}}$ with a displacement feature $\left\lvert\hat{\mathbf{D}}_{\mathbf{p}^{k}_{i}}-z_{\mathbf{p}^{k}_{i}}\right\rvert$ that compares the rendered depth for input views and the depth $z_{\mathbf{p}^{k}_{i}}$ of a primitive, allowing for occlusion-aware reasoning.
@@
$\text{SH}_{i,k}^{\textit{residuals}}$ $=\text{MLP}\left(\text{CrossAttn}\left(\mathcal{X}_{\mathbf{p}^{k}_{i}},\mathbf{V}^{i}_{\mathcal{G}}\right)\right)\text{,}$ (6)
@@
Furthermore, both coarse and fine modules are differentiable and updated simultaneously. Thus, the fine renderings can further regularize the coarse predictions.
@@ 4 Implementation Details
Datasets.   We train our model on multi-view synthetic renderings of objects [70, 49], based on the Objaverse dataset [15], which includes 264,775 scenes with a train/test split of 10:1. Each scene contains 38 circular views with an image resolution of $512\times 512$. To ensure sufficient angular coverage of the input views, we employ the classical K-means algorithm to cluster the cameras into 4 clusters. We employ eight views for supervision for each object and leverage the loss objectives outlined in Eq. 8 to update the network.
@@
Network.   We developed LaRa using PyTorch Lightning [20] and conducted our training on 4 NVIDIA A100-40G GPUs over a period of 2 days for the fast model and 3.5 days for the base model, with a batch size of 2 per GPU. We use DINO-base for encoding $M\!=\!4$ multi-view images at a resolution of $512\times 512$. We use a volume resolution of $W\!=\!16$ with $C\!=\!768$ channels for the image feature volume, and a resolution of $W\!=\!32$ with $C\!=\!256$ channels for the embedding volume, dividing both into $G\!=\!16$ groups for the group attention layers. Our group attention network consists of 12 layers, producing a Gaussian volume of size $64\!\times\!64\!\times\!64\!\times\!80$. We choose $K\!=\!2$ primitives for each voxel, and constrain the offset radius to $r=1/32$ of the length of the bounding box. The total number of trainable parameters is 125 million.
@@ 5.1 Comparison
As shown in Table 1, our approach is robust to scene scale and can be generalized to real captured images, such as those in the Co3D dataset, thanks to our canonical modeling and projection-based feature lifting. In contrast, LGM [60] leverages a monocular prediction and fusion technique that requires a reference scene scale and a constant camera-object distance to avoid focal length and distance ambiguity. This requirement significantly limits its generalizability to real data. As shown in Table 1 and  Figure 5, LGM provides faithful reconstructions in datasets with a strict constant camera-object distance, such as GSO, but fails to generalize to unconstraint multi-view data such as in Objaverse and Co3D datasets, and exhibits serious distortions. Our model trained on 4 A100-40G GPUs for 2 days demonstrates superior results compared to the LGM model trained on 32 A100-80G GPUs (8$\times$ GPUs, 16$\times$ RAM, 32$\times$ GPU hours) and on the same synthetic Objaverse dataset [15].
@@
Geometry.   We evaluate the quality of our geometry reconstruction by comparing the depth reconstructions on novel views, generated by a weighted sum of the $z$ values of the primitives. As shown in Table 2, our approach achieves significantly lower L1 errors and higher geometry accuracy than other baselines. In Figure 6, we also visualize geometry reconstruction by extracting meshes using TSDF. In addition, our trajectory video rendering (48 views at a resolution of $512$) together with mesh extraction is highly efficient, as it does not require fine-tuning and can be performed in just 2 seconds.
@@ 5.2 Ablation Study
We now analyze the contributions of individual elements of our model design.
@@ 6 Conclusion
We have presented LaRa, a novel method for reconstructing 360 bounded radiance fields given large-baseline inputs. Our central idea is to match image features and embedding volume through unified local and global attention layers. By integrating this with a coarse-fine decoding and splatting process, we achieve high efficiency for both training and inference. In future work, we plan to explore how to enlarge the batch size per-GPU and volume resolution without increasing GPU usage.
Appendix 0.A Limitations and Discussions
Our LaRa demonstrates a remarkable efficiency feed-forward model that achieved high-fidelity all-around novel-view synthesis and surface reconstruction from sparse large-baseline images. However, our approach struggles to recover high-frequency geometry and texture details, mainly due to the low volume resolution. Enhancing our approach with techniques such as gradient checkpointing or mixed-precision training can potentially increase training batch size as well as volume resolution. We have also noticed that our method can yield inconsistent rendering results when the geometry is incorrectly estimated or when reconstructing multi-view inconsistent inputs, as demonstrated in the comparison video. This occurs because our method utilizes second-order Spherical Harmonic appearance modeling. Although such modeling can capture view-dependent effects, it also introduces a stronger ambiguity between geometry and appearance. We believe that incorporating our method with a physically-based rendering process can potentially address this issue. In addition, our work assumes posed inputs, but estimating precise camera poses for sparse views is a challenge in practice. Incorporating a pose estimation module [66] into the feed-forward setting is an orthogonal direction to our work.
Appendix 0.B Acknowledgements
We thank Bozidar Antic for pointing out a bug that resulted in an improvement of about 1dB. Special thanks to BinBin Huang and Zehao Yu for their helpful discussion and suggestions. We would like to thank Bi Sai, Jiahao Li, Zexiang Xu for providing us with the testing examples of Instant3D, and Jiaxiang Tang for helping us to construct a comparison with LGM. This project was supported by the ERC Starting Grant LEGO-3D (850533) and the DFG EXC number 2064/1 - project number 390727645.
Appendix 0.C More Visual Results
Finally, we report more qualitative results on the testing sets, as shown in Figure 8,Figure 9, and Figure 10.
Figure 8: Reconstruction results on Gobjaverse testing set.
Figure 9: Reconstruction results on Google Scanned Object dataset.
Figure 10: Reconstruction results on Instant3D scenes.